A 0.0625 kg ingot of metal is heated to 213◦C
and then is dropped into a beaker containing
0.4 kg of water initially at 20◦C.
If the final equilibrium state of the mixed
system is 22.4◦C, find the specific heat of
the metal. The specific heat of water is
4186 J/kg ·◦ C.
Answer in units of J/kg ·◦ C.
1
Expert's answer
2020-04-06T08:55:04-0400
"mc(T_m-T)=Mc_w(T-T_w)"

"(0.0625)(213-22.4)c=(0.4)(4186)(22.4-20)"

"c=337\frac{J}{kg \degree C}"